f(x)=x2 from x=0 to x=5


1
Expert's answer
2022-05-03T11:33:30-0400
"Area=\displaystyle\int_{0}^{5}x^2dx=[\dfrac{x^3}{3}]\begin{matrix}
 5 \\
 0
\end{matrix}=\dfrac{125}{3}({units}^2)"